# Mitigating Information Loss in Tree-Based Reinforcement Learning via Direct Optimization

### v3

@@ Abstract
Reinforcement learning (RL) has seen significant success across various domains, but its adoption is often limited by the black-box nature of neural network policies, making them difficult to interpret. In contrast, symbolic policies allow representing decision-making strategies in a compact and interpretable way. However, learning symbolic policies directly within on-policy methods remains challenging. In this paper, we introduce SYMPOL, a novel method for SYMbolic tree-based on-POLicy RL. SYMPOL employs a tree-based model integrated with a policy gradient method, enabling the agent to learn and adapt its actions while maintaining a high level of interpretability. We evaluate SYMPOL on a set of benchmark RL tasks, demonstrating its superiority over alternative tree-based RL approaches in terms of performance and interpretability. Unlike existing methods, it enables gradient-based, end-to-end learning of interpretable, axis-aligned decision trees within standard on-policy RL algorithms. Therefore, SYMPOL can become the foundation for a new class of interpretable RL based on decision trees.
@@ 2 Related Work
Recently, the integration of symbolic methods into RL has gained significant attention. Symbolic RL does cover different approaches including program synthesis (Trivedi et al., 2021; Penkov & Ramamoorthy, 2019; Verma et al., 2018), concept bottleneck models (Ye et al., 2024), piecewise linear networks (Wabartha & Pineau, 2024), logic (Delfosse et al., 2024b), mathematical expressions (Landajuela et al., 2021; Guo et al., 2024; Luo et al., 2024; Xu et al., 2022). Another line of work aims to synthesize symbolic policies using logical rules, leveraging differentiable inductive logic programming for gradient-based optimization (Jiang & Luo, 2019; Cao et al., 2022). In contrast to first-order rules, DTs offer greater flexibility by not only combining atomic conditions but also comparing features against thresholds — a critical capability for handling continuous observation spaces. In this paper, we focus exclusively on tree-based methods for symbolic RL. Several approaches have been proposed to leverage the strengths of interpretable, tree-based representations within RL frameworks. However, each approach comes with its own critical limitations. We summarize existing methods into three streams of work:
@@
(1) Post-processing.   One line learns full-complexity policies first and then performs some kind of post-processing for interpretability. One prominent example is the VIPER algorithm (Bastani et al., 2018). In this case, a policy is learned using NNs before DTs are distilled from the policy. However, distillation methods often suffer from significant performance mismatches between the training and evaluation policies (see Figure 1(b)). To mitigate this mismatch, existing methods often learn large DTs (VIPER learns DTs with 1,000 nodes) and therefore aim for systematic verification rather than interpretability. In contrast, SYMPOL is able to learn small, interpretable DTs (average of only 50 nodes) without information loss. Following VIPER, various authors proposed similar distillation methods (Li et al., 2021; Liu et al., 2019; 2023; Jhunjhunwala et al., 2020). Furthermore, Kohler et al. (2024) propose a novel distillation method that distills interpretable and editable programmatic tree policies. In contrast to SYMPOL, the extracted trees are not considered axis-aligned, as they allow for linear combinations and multiple features within the internal nodes.
@@
(2) Custom optimization.   The third line involves custom, tree-specific optimization techniques and/or objectives (Ernst et al., 2005; Roth et al., 2019; Gupta et al., 2015; Kanamori et al., 2022) and, hence, is more time-consuming and less flexible. As a result, their policy models cannot be easily integrated into existing learning RL frameworks. Examples are evolutionary methods (Costa et al., 2024; Custode & Iacca, 2023) and linear integer programming (Vos & Verwer, 2023). Topin et al. (2021) propose Iterative Bounding Markov Decision Process (IBMDP) that allow learning DT policies through a masking procedure and modified value updates by using arbitrary function approximators. However, using IBMDP, the learning problem becomes more complex compared to the base MDP, which can result in poor scalability and limits the applicability to very simple tasks (Milani et al., 2022; Kohler et al., 2024). In contrast, SYMPOL optimizes a DT policy directly on the base MDP, avoiding these limitations.
@@
Furthermore, trees have also been in used in other agentic components than the policy, such as reward functions (Milani et al., 2022; Kalra & Brown, 2023; 2022). Similarly, ensemble methods (Fuhrer et al., 2024; Min & Elliott, 2022) have been proposed. However, policies consisting of hundreds of trees and nodes lack interpretability and therefore are out of scope for this paper.
@@ 5.1 Experimental Settings
Setup.   We implemented SYMPOL in a highly efficient single-file JAX implementation that allows a flawless integration with highly optimized training frameworks (Lange, 2022; Weng et al., 2022; Bonnet et al., 2024). Our implementation is available in the supplementary material, and will be made publically available upon acceptance. We evaluated our method on several environments commonly used for benchmarking RL methods. Specifically, we used control environments including CartPole (CP), Acrobot (AB), LunarLander (LL), MountainCarContinuous (MC-C) and Pendulum (PD-C), as well as the MiniGrid (Chevalier-Boisvert et al., 2023) environments Empty-Random (E-R), DoorKey (DK), LavaGap (LG) and DistShift (DS).

### v4

@@ Abstract
Reinforcement learning (RL) has seen significant success across various domains, but its adoption is often limited by the black-box nature of neural network policies, making them difficult to interpret. In contrast, symbolic policies allow representing decision-making strategies in a compact and interpretable way. However, learning symbolic policies directly within on-policy methods remains challenging. In this paper, we introduce SYMPOL, a novel method for SYMbolic tree-based on-POLicy RL. SYMPOL employs a tree-based model integrated with a policy gradient method, enabling the agent to learn and adapt its actions while maintaining a high level of interpretability. We evaluate SYMPOL on a set of benchmark RL tasks, demonstrating its superiority over alternative tree-based RL approaches in terms of performance and interpretability. Unlike existing methods, it enables gradient-based, end-to-end learning of interpretable, axis-aligned decision trees within standard on-policy RL algorithms. Therefore, SYMPOL can become the foundation for a new class of interpretable RL based on decision trees. Our implementation is available under: https://github.com/s-marton/sympol
@@ 2 Related Work
Recently, the integration of symbolic methods into RL has gained significant attention. Symbolic RL does cover different approaches including program synthesis (Trivedi et al., 2021; Penkov & Ramamoorthy, 2019; Verma et al., 2018), concept bottleneck models (Ye et al., 2024), piecewise linear networks (Wabartha & Pineau, 2024) and mathematical expressions (Landajuela et al., 2021; Guo et al., 2024; Luo et al., 2024; Xu et al., 2022). Another line of work aims to synthesize symbolic policies using logical rules, leveraging for instance differentiable inductive logic programming for gradient-based optimization (Delfosse et al., 2024b; Jiang & Luo, 2019; Cao et al., 2022). In contrast to first-order rules, DTs offer greater flexibility by not only combining atomic conditions but also comparing features against thresholds — a critical capability for handling continuous observation spaces. Furthermore, trees have also been in used in other agentic components than the policy, such as reward functions (Milani et al., 2022; Kalra & Brown, 2023; 2022). In this paper, we focus exclusively on tree-based methods for symbolic RL. Similarly, ensemble methods (Fuhrer et al., 2024; Min & Elliott, 2022) have been proposed. However, policies consisting of hundreds of trees and nodes lack interpretability and therefore are out of scope for this paper. Several approaches have been proposed to leverage the strengths of interpretable, tree-based representations within RL frameworks, each approach comes with its own critical limitations. We summarize existing methods into three streams of work:
@@
(1) Post-processing.   One line learns full-complexity policies first and then performs some kind of post-processing for interpretability. One prominent example is the VIPER algorithm (Bastani et al., 2018) where a NN policy is learned before distilling a DT from the policy. However, distillation methods often suffer from significant performance mismatches between the training and evaluation policies (Figure 1(b)). To mitigate this mismatch, existing methods often learn large DTs (VIPER learns DTs with 1,000 nodes) and therefore aim for systematic verification rather than interpretability. In contrast, SYMPOL is able to learn small, interpretable DTs (average of only 50 nodes) without information loss. Following VIPER, various authors proposed similar distillation methods (Li et al., 2021; Liu et al., 2019; 2023; Jhunjhunwala et al., 2020). Furthermore, Kohler et al. (2024) propose a novel method that distills interpretable and editable programmatic tree policies. In contrast to SYMPOL, the extracted trees are not considered axis-aligned, as they allow for linear combinations and multiple features within the internal nodes.
@@
(2) Custom optimization.   Methods involving custom, tree-specific optimization techniques and/or objectives (Ernst et al., 2005; Roth et al., 2019; Gupta et al., 2015; Kanamori et al., 2022) are generally more time-consuming and less flexible. As a result, their policy models cannot be easily integrated into existing learning RL frameworks. Examples are evolutionary methods (Costa et al., 2024; Custode & Iacca, 2023) and linear integer programming (Vos & Verwer, 2023). Topin et al. (2021) propose Iterative Bounding Markov Decision Process (IBMDP) that allow learning DT policies through a masking procedure and modified value updates by using arbitrary function approximators. However, using IBMDP, the learning problem becomes more complex compared to the base MDP, which can result in poor scalability and limits the applicability to very simple tasks (Milani et al., 2022; Kohler et al., 2024). In contrast, SYMPOL optimizes a DT policy directly on the base MDP, avoiding these limitations.
@@ 5.1 Experimental Settings
Setup.   We implemented SYMPOL in a highly efficient single-file JAX implementation that allows a flawless integration with highly optimized training frameworks (Lange, 2022; Weng et al., 2022; Bonnet et al., 2024). [^1] Our implementation is available under https://github.com/s-marton/sympol. We evaluated our method on several environments commonly used for benchmarking RL methods. Specifically, we used the control environments CartPole (CP), Acrobot (AB), LunarLander (LL), MountainCarContinuous (MC-C) and Pendulum (PD-C), as well as the MiniGrid (Chevalier-Boisvert et al., 2023) environments Empty-Random (E-R), DoorKey (DK), LavaGap (LG) and DistShift (DS).